a bag contain 3 red balls,4 yellow balls,5 blue balls.find the probability of selecting at random a red ball.
Answer by jim_thompson5910(28595)

Notation: P(Red) denotes the probability of selecting a red ball


First, add up the number of balls to get a total of 3+4+5=12 balls total.


P(Red) = Number of Red Balls/Number of Balls Total = 3/12 = 1/4


So the probability of selecting a red ball is 1/4 which is 0.25 in decimal form giving us a 25% chance.